you are deciding between two cars with different engines and want the bigger of the two. One engine displaces 350 cubic inches. The other displaces 5,500 cubic centimeters. By how many cubic centimeters is the bigger engine larger than the smaller one?
Answer by checkley77(12844)

1 in. = 2.54 cm.
cubert350=7.0473 in.
7.0473*2.54=17.9 cm.
17.9^3=5,735 cm^3 is the 350 in^3 engine size.
5,735-5,500=235 cm^3.
the 350in^ engine is 235 cm^3 larger.